OpenStudy (anonymous):

find the sum of the given arithmetic series. (: PLEASE HELP ME!!! 1.30+23+16+9+...(-30) 2. 26+19+12+5+...+(-37) 3.27+20+13+6+...+(-36)

OpenStudy (anonymous):

just use a caclulator lol

OpenStudy (anonymous):

and it will give me the answers? lol

jagr2713 (jagr2713):

dont listen to @mitu12 Mathway is a site that gives the answers and doesnt explain. it does but you have to pay and sometimes you dont understand the concept.

OpenStudy (anonymous):

ohhh okay thank you!

OpenStudy (mitu12):

its like a calculator

zepdrix (zepdrix):

The sum of an arithmetic series is given by:\[\Large\rm S_n=\frac{n(a_1+a_n)}{2}\]

zepdrix (zepdrix):

We're given the last term and the first term. The work we'll have to do is figuring out `how many` terms we have. That is our \(\Large\rm n\) value.

zepdrix (zepdrix):

The first is \(\Large\rm a_1=30\) The common difference is \(\Large\rm d=a_2-a_1=-7\) The general form for finding the nth term of a geometric sequence is given by:\[\Large\rm a_n=a_1+d(n-1)\]Plugging in our stuff gives us:\[\Large\rm a_n=30-7(n-1)\]

zepdrix (zepdrix):

We know that our LAST term is -30. Let's plug that in for our \(\Large\rm a_n\) and see which \(\Large\rm n\) value that corresponds to. That will tell us how many terms we have.

zepdrix (zepdrix):

\[\Large\rm -30=30-7(n-1)\]

zepdrix (zepdrix):

Do you understand how to solve for n in that set up?

OpenStudy (anonymous):

do i just try and get n by itself?

zepdrix (zepdrix):

Yes, then we will have an \(\Large\rm n\) value for our \(\Large\rm S\) formula :)

OpenStudy (anonymous):

okay!

zepdrix (zepdrix):

Hmm it's not working out to an integer value.. this is really weird :o

zepdrix (zepdrix):

Oh I see the problem. The first one listed is NOT an arithmetic series. One of the numbers is messed up. We cant have 30 AND -30 in the series. Did you copy it down correctly? I assume you just copy/pasted, yes? :o

OpenStudy (anonymous):

sorry i meant (-33)

zepdrix (zepdrix):

Oh ok that would make more sense :3 So we're actually plugging it in like this:\[\Large\rm -33=30-7(n-1)\]

OpenStudy (anonymous):

okay sorry about that again!!

OpenStudy (anonymous):

-55n?

zepdrix (zepdrix):

Hmmm, I'm not sure what you did there. We're trying to isolate our n.\[\Large\rm -33=30-7(n-1)\]So start by subtracting 30 from each side,\[\Large\rm -63=-7(n-1)\]Divide by -7,\[\Large\rm 9=n-1\]Add 1 to each side,\[\Large\rm 10=n\]Understand those steps ok? :o

OpenStudy (anonymous):

yes. lol

zepdrix (zepdrix):

So now we have all of the information we need: \(\Large\rm a_1=30, \qquad a_n=-33,\qquad n=10\) Plug them into your formula for sum,\[\Large\rm S=\frac{n(a_1+a_n)}{2}\]What do you get for your sum? :)

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

do you mind helping me with the other two? it is okay if you do not want too.

zepdrix (zepdrix):

So you'll need to do these steps for all of them, so get comfortable with the process. \[\Large\rm S=\frac{n(a_1+a_n)}{2}\]We need three things, \(\Large\rm a_1,~ a_n,~n\) We already have \(\Large\rm a_1,~a_n\) In order to find \(\Large\rm n\), we have to first find the common difference and plug everything into our formula for the nth term.\[\Large\rm a_n=a_1+d(n-1)\]

zepdrix (zepdrix):

So start by finding your common difference. (The second term minus the first term). Then plug all that jazz into the formula to find your n.

OpenStudy (anonymous):

okay thank you!!

zepdrix (zepdrix):

You can do it mads! c: Lemme know if you get stuck. Look back at the steps we took also.

OpenStudy (anonymous):

was the -15 correct? lol

zepdrix (zepdrix):

For the first one? Yes, good job c:

OpenStudy (anonymous):

Thank you for all your help(:

OpenStudy (anonymous):

i got -11 for the 2nd one and -9 for the 3rd. Do you mind checking to see if i got the correct answer? :)

zepdrix (zepdrix):

For the second problem, \[\Large\rm a_1+a_n=-11\]Did you forgot to multiply that by \(\Large\rm \dfrac{n}{2}\)? Remember,\[\Large\rm S=\frac{n}{2}(a_1+a_n)\]

zepdrix (zepdrix):

You have to find your n value :O

OpenStudy (anonymous):

ive confused myself now i am lost..im so sorry.

OpenStudy (anonymous):

i got,-10.57 but i think i did something wrong on my work.

zepdrix (zepdrix):

What is the common difference? \(\Large\rm d=?\)

OpenStudy (anonymous):

-7

zepdrix (zepdrix):

\[\Large\rm a_n=a_1+d(n-1)\]\[\Large\rm -37=26-7(n-1)\]And what'd you get for n? :d

OpenStudy (anonymous):

-10.57?

zepdrix (zepdrix):

Hmm, no. Make sure you do the subtraction step before you do the division step.

OpenStudy (anonymous):

okay, i see what i did wrong...

zepdrix (zepdrix):

That's only half of the work though! Remember you have to take your n value and plug it into the S formula.

OpenStudy (anonymous):

nevermind i messed up again. i am so sorry.

zepdrix (zepdrix):

lol :) dat mads

zepdrix (zepdrix):

Is this step ok?